kwitonda:

. A 50.00-mL sample containing an analyte gives a signal of 11.5 (arbitrary units). A second 50-mL aliquot of the sample, which is spiked with 1.00-mL of a 10.0-ppm standard solution of the analyte, gives a signal of 23.1. What is the concentration of analyte in the original sample?

Vocaloid:

Not 100% sure about this, but this is my approach: A=εbC where ε is the molar absorptivity constant (since both samples contain the same analyte, assume ε is constant for both), b is the path length (assuming you're measuring both samples using the same size cell, assume b is the same for both samples), and C is concentration letting the first 50 mL sample have absorbances A1 and C1, and the spiked sample having absorbances A2 and C2 C1 = x ppm = x mg / (50/1000 L) meaning the mass x = C1 * (50/1000 L) now, looking at solution 2 C2 = (the mass of a 50 mL aliquot + the mass of the 1.0 mL spike) / (50 + 1 mL) or converting to the more appropriate ppm (mg/L) units C2 = (C1 * (50/1000 L) + (10.0mg/L) * (1/1000)L) / (51/1000 L) which gives us C2 in terms of C1, in ppm going back to the original Beer's equation: A1=εbC1 A2=εbC2 dividing these gives us A1/A2 = C1/C2 you have C2 in terms of C1, so you can plug in the C2 equation, and the absorbances, and solve for C1